RC meeting and Run Plan, Wednesday December 7th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Daily at 3:30pm

  • Location: Counting house conference room

Shift Schedule

Accelerator Status

  • RF recovery day on Wednesday day 8:00AM - 4:00PM

Current Status

  • The Hall will remain controlled access till 4 PM today
  • There's a long beam shutdown yesterday due to accelerator RF issues, after the shutdown, per Bill's suggestion, we did a harp can to make sure the beam spot size on our target is still good. Thanks to Jiwan for performing the beam restoration task! One thing during the beam restoration yesterday swing is that the MCC found our Moller quads were on for the past three weeks, which are supposed to be off. After a few unsuccessful trials delivering our desired beam spot size, MCC turned off Moller quads ( https://logbooks.jlab.org/entry/4097518 ). After that, MCC still have troubles delivering our desired beam spot size (320 - 350 on target), We talked with Yves, Gordon, Arun, Bill, and Simona and agreed that we should accept the current result ( https://logbooks.jlab.org/entry/4097607 ). Please refer to a more detailed log from Iuliia ( https://logbooks.jlab.org/entry/4097400 )
  • After beam restoration, we did a carbon hole check, we found the hole in the right position with the previous beam position, no scan was needed, see the carbon hole ( https://logbooks.jlab.org/entry/4097641 )
  • Resumed data taking on He3 target, 36 uA Beam, in the owl shift

Issues Resolved

  • hodoscope fixed
  • Hunter finished Target calibration
  • APV remain the same

Issues remaining

  • beam step size jump is still unresolved

Short Term Run Plan

  • Continue production run on He3 target, 36 uA
  • Use GEnII-noSBSGEMs DAQ config
  • Use ps3=0, ps5=6, others =-1
  • BBCal threshold at -366 mV, HCal threshold at -19 mV
  • NMR measurement every 3 - 4 hours, close to 4 hours is better

Planned Accesses to the Hall

  • Access for hodoscope
  • Access for GEM BB APV will need a further evaluation see if it's necessary

Special Requests

  • BigBite GEMs might need new alignment, since we worked on it this morning

Notices in Effect for Shift Crew

  • Power down SBS, BB, and Detector HV before making an access to the Hall

Data Collection Status

  • 73 out of 141 days (51.7%)
  • Uncorrected accumulated charge on He3: 45.588 C out of 232C (19.65%)